OpenStudy (anonymous):

Which of the following fractions can be used in the conversion of 32 m3 to the unit mm3? fraction cubed: one mm over ten to the negative third power m fraction cubed: one m over ten to the negative third power mm one mm over ten to the negative third power m ten to the negative third power m over one mm

OpenStudy (festinger):

When converting units i like to use the method which is often applied to fractions, which is to "cancel out" the terms. \[1mm=10^{-3}m\]I choose to rearrange this so that I can cancel off the units. In this case to cancel off the mm. \[\frac{10^{-3}m}{1mm}=1\] So\[32mm=32mm*1= 32mm*\frac{10^-3m}{1mm}=0.032m\] In your question, it is: So\[32mm^{3}=32mm^{3}*1= 32mm^{3}*(\frac{10^-3m}{1mm})^{3}=32mm^{3}*(\frac{1m}{1000mm})^{3}\]

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

Can you help me with another one?

OpenStudy (anonymous):

Which of the following is not a step in the conversion of 1.3 dam3 to L? fraction cubed: ten to the negative third power L over one mL one mL over one cubic cm fraction cubed: one cm over ten to the negative second power m fraction

OpenStudy (festinger):

Is there a fourth option? The 3 look fine to me.

OpenStudy (anonymous):

yes

OpenStudy (festinger):

fraction cubed: ten to the negative third power L over one mL I think this answer appears like\[(\frac{10^{-3}L}{1mL})^{3}\]